Solve the equation for the interval [0, 2π).
cos^2x + 2 cos x + 1 = 0   

2 sin^2x = sin x

cos x = sin x

sec^2x - 2 = tan^2x
Answer #1

Cosx = COST x = X cos tacosxt1=0 cose + cosa + cosx + 1 =0 COLIC Code + 1) + 1 (cosse +1)=0 Cosx + 1)² = 0 COOK+1=0 Cosa = -13 lan Cosa = Susa COS Cosa tanx = tana tas (774) x = 7/4 is tve is quadrant D, atx AT 7/4 4THT = ST 4 secx - 2 = tantx It tan